OpenStudy (callisto):

the position vectors of P, Q and R are i+5j+k, -i-2j-2k and -j-k respectively. find ∠PQR. How should i start ?

OpenStudy (experimentx):

PQ = Q-P QR = R-Q |PQ||QP|cos(theta) = PQ.dot.QP

OpenStudy (ash2326):

 Find Vector P-Q And Vector R-Q Find the angle by using \[(P-Q).(R-Q)=|P-Q||R-Q| \cos \theta\] Find theta using above. Can you do it?

OpenStudy (callisto):

let me try, thanks

OpenStudy (callisto):

Is it PQ.RQ or QP.QR? and why?

OpenStudy (ash2326):

P, Q and R are position vectors of point P, Q and R. If we want a vector directed from Q to P then it's P-Q or PQ. and a vector directed from Q to R then it's R-Q or RQ either choose PQ and RQ to find angle or choose QP and QR to find the angle

OpenStudy (anonymous):

1)Find vector PQ and QR 2) Use vector Dot " . " product for angle, > > PQ= -2i -7j-3k and QR= i+j+k > now PQ . QR = |PQ| |QR| Cos (theta) => 12 = (62)^1/2 (3)^1/2 Cos (theta)

OpenStudy (callisto):

why is that when a vector is directed from Q to P, it is PQ but a vector directed from Q to R, it is RQ

OpenStudy (anonymous):

technically its not 'PQ', thing IS we'r lazy, and wont spend a sec on something that basic

OpenStudy (anonymous):

Sorry, but how do I mention a person in post ? ? ? my browser seems unresponsive . . .

OpenStudy (callisto):

i understand that QR is from Q to R (alright i can't type the arrow signs here, just assume there are arrows) but why isn't PQ from P to Q? to mention a person in post just type @ followed by the name of that person

OpenStudy (anonymous):

 @Callisto PQ is indeed directed from P to Q, similarly a vec. from Q to P would be QP